A continuous cable of total length 4 m is wrapped around the small pulleys at A, B, C, and D. If each spring is stretched 340use k 500N/M and 250mm instead of 340mm
Answer #1

FBD FBD os pulley A de Fspring=F= F = kxsl ol=250 mm= 0.25 m K = 500nm F=F<=125N (500x0.250) = 125N A E=125n FBD of pulley C150- 25 ain o = 0 75 T = - Sino Stoeched-FBD 0 Oiasm CC,=> deflection of Spring a A A, =) deflection of spoing A CC=AA,=0-25mSubstitute, 0= 48.59 in eqD 75 TOON sin (48.59) Consider FBD (3) ► +ve -mg+ at coso = 0 21 COSO 2 100 X COS (48-59) 9.81 3 in